Given that V=r²p/4hl,where p/l=pressure gradient,h=coefficient of viscosity,obtain an expression for the dimensionless constant in the expression V
1
Expert's answer
2019-01-21T10:53:52-0500

All variables have units:

"\frac{p}{l}=[\frac{\text{Pa}}{\text{m}}],""r=[\text{m}],""h=[\text{Pa}\cdot \text{s}],""v=\frac{r^2p}{4hl}=\frac{[\text{m}^2]\cdot [\text{Pa}]}{4[\text{Pa}\cdot \text{s}]\cdot[\text{m}]}=[\frac{\text{m}}{\text{s}}]."

For the dimensionless constant we have

"4=\frac{r^2p}{vhl}."